ganeshie8 (ganeshie8):

A) How many ordered sets of non-negative integers \((x, y)\) are there such that \(x+y \le 20\) ? B) generalize the result to \(x+y \le n\)

OpenStudy (thomas5267):

I think A) is 21 and B) is n+1. Don't count on me though.

ganeshie8 (ganeshie8):

thats right if it is equality : x+y = 20

OpenStudy (thomas5267):

I have a idea! We can use recurrence relations to find this out. Let \(N(eq)\) be the function that returns the number of ordered set that satisfies the equation \(eq\). For \(n=0\), \(N(x+y\leq0)=1\) For \(n=1\), \(N(x+y\leq1)=3=N(x+y\leq0)+N(x+y=1)\) For \(n=2\), \(N(x+y\leq2)=N(x+y\leq1)+N(x+y=2)\) Did I do something wrong?

ganeshie8 (ganeshie8):

looks perfect to me ! you want to iterate n=0..20 is it ?

ganeshie8 (ganeshie8):

F(n) = F(n) + F(n-1)

OpenStudy (thomas5267):

Ah. Let \(E(n)=N(x+y=n)\). \(E(n)=n+1\). \[ \begin{align*} N(x+y\leq n)&=\sum_{k=0}^nE(k)\\ &=\sum_{k=0}^n(k+1)\\ &=\frac{n(n+1)}{2}+n+1\\ &=\frac{(n+1)(n+2)}{2} \end{align*} \]

ganeshie8 (ganeshie8):

Exactly ! that expression looks pretty xD n=0 : (0+1) n=1 : (1+1) n=2 : (2+1) ... n=n : (n+1) ------------- n(n+1)/2 + (n+1)

ganeshie8 (ganeshie8):

i think this should work : Number of non-negative integer solutions for \(x+y \le n\) is\[\large \binom{n+1}{2}\]

ganeshie8 (ganeshie8):

may be we can go ahead and generaluze this to \(\large ax+by \le n\) xD not sure whether it would be easy..

OpenStudy (thomas5267):

Have to get my dinner now. Chat later.